Bond A and bond B have par values of $1,000. They both have coupon rates of...

  1. Bond A and bond B have par values of $1,000. They both have coupon rates of 10%, and they both pay semi-annual interest. Bond A has 3 years left to maturity and bond B has 14 years left to maturity.
  1. Calculate the value of each bond for a required return of 5%, 10%, and 15%.

Required return

Value bond A

Value bond B

5%

?

?

10%

?

?

15%

?

?

  1. Which bond is more responsive to a change in required return? Why?
Answer #1

1 Required Return value of Bond A Value of Bond B 5% 1137.70 1499.12 1000.00 1000.00 15% 882.65 710.67 10% w

C B 1 Required Return value of Bond A Value of Bond B 2 0.05 =PV(A2/2,2*3,-10%*1000/2,-1000) =PV(A2/2,2*14,-10%*1000/2,-1000)

Bond B is more sensitive to change in required return because of higher maturity and sensitivity increases with maturity